I am working on homework we are doing graphing with slopes Not sure how to do this part of the homework the question is y+x=3; (1,) (3,),(2,) it says Graph the equation by determining the missing values needed to plot the ordered pairs.
Answer by Alan3354(69443)

I am working on homework we are doing graphing with slopes Not sure how to do this part of the homework the question is y+x=3; (1,) (3,),(2,) it says Graph the equation by determining the missing values needed to plot the ordered pairs.
-----------
y+x=3 --> y = 3-x
=========
Sub the values for x and find y.
Plot the points.
Draw a line thru them.